### v2

# Selection-Induced Contraction of Innovation Statistics in Gated Kalman Filters

Abstract
Validation gating is a fundamental component of classical Kalman-based tracking systems. Only measurements whose normalized innovation squared (NIS) falls below a prescribed threshold are considered for state update. While this procedure is statistically motivated by the chi-square distribution, it implicitly replaces the unconditional innovation process with a conditionally observed one, restricted to the validation event.
This paper shows that innovation statistics computed after gating converge to gate-conditioned rather than unconditional nominal reference quantities. Under classical linear-Gaussian assumptions, we derive exact expressions for the first- and second-order moments of the innovation conditioned on ellipsoidal gating, and show that gating induces a deterministic, dimension-dependent contraction of the innovation covariance relative to the nominal reference.
The analysis is extended to nearest-neighbor (NN) association, which is shown to act as an additional statistical selection operator. We prove that selecting the minimum norm innovation among multiple in-gate measurements introduces an unavoidable energy contraction, implying that nominal innovation reference statistics cannot be preserved under nontrivial gating and association due to deterministic selection effects, even under perfectly matched linear-Gaussian Kalman filter assumptions. Closed-form results in the two-dimensional case quantify the combined effects and illustrate their practical significance.
I Introduction
Kalman-based tracking systems are a cornerstone of modern aerospace, radar, and navigation applications, where reliable state estimation is required under uncertainty. A central role in such systems is played by innovation-based statistics, which are used for measurement validation, data association, consistency monitoring, and adaptive tuning. Among these, the normalized innovation squared (NIS) is particularly attractive due to its simple statistical characterization under nominal linear Gaussian assumptions: in the absence of additional selection mechanisms, the NIS follows a chi-square distribution whose moments provide natural reference values for diagnostic and tuning procedures.
In operational tracking systems, however, innovation statistics are rarely observed in their unconditional form. Prior to data association and state update, measurements are typically subjected to ellipsoidal validation gating, whereby only innovations whose Mahalanobis distance falls below a prescribed threshold are accepted. Validation gating is widely justified on statistical and practical grounds, as it limits computational complexity and suppresses gross outliers. As a result, nearly all innovation-based diagnostics used in practice operate on a post-gate innovation stream rather than on the nominal innovation process assumed in classical Kalman filter theory.
Despite its ubiquity, the statistical consequences of validation gating are rarely treated explicitly. Once gating is applied, the innovations that enter association, filtering, and diagnostic logic are no longer samples from the nominal Gaussian distribution, but from a distribution truncated to the validation region. Consequently, innovation-based statistics computed after gating need not satisfy the classical reference properties associated with the chi-square law. In particular, empirical NIS statistics obtained from accepted measurements may systematically deviate from their nominal expectations even when the underlying system model and noise statistics are perfectly matched.
A closely related issue arises in data association. When multiple measurements fall inside the validation gate, NN association is commonly employed to select a single measurement for the state update. While computationally efficient and widely used, NN association further conditions the observed innovation through order-statistic selection, as it selects the innovation with minimum normalized distance among multiple candidates. The combined effect of validation gating and NN association therefore induces a multi-stage selection process whose impact on innovation statistics is not captured by nominal Kalman filter assumptions.
The objective of this paper is to isolate and characterize the statistical effects induced solely by validation gating and NN association, independent of clutter, false alarms, nonlinear dynamics, or modeling errors. Within the classical linear–Gaussian Kalman filtering framework, we provide an exact characterization of the innovation moments conditioned on ellipsoidal gating and show that validation gating induces a deterministic, dimension-dependent contraction of the innovation covariance relative to the nominal reference. We further show that NN association acts as an additional statistical selection operator and introduces an unavoidable, multiplicity-dependent contraction of innovation energy through order-statistic selection. Together, these results imply that nominal innovation statistics cannot be preserved under nontrivial gating and association, even when all Kalman filter assumptions are satisfied.
I-A Related Work
Innovation-based statistics are a fundamental component of Kalman filtering theory and practice, where quantities such as the NIS are routinely used for consistency monitoring, measurement validation, and adaptive tuning. Under nominal linear Gaussian assumptions, the statistical properties of the innovation process are well understood and documented in classical estimation references [10, 9, 8, 3, 4].
In practical tracking systems, innovation statistics are commonly employed in adaptive filtering and noise covariance estimation schemes, often relying on assumed chi-square properties of the NIS [12, 1, 20, 21]. Such approaches implicitly assume that the observed innovation sequence is representative of the nominal innovation distribution. However, in operational settings, innovation statistics are almost always evaluated after validation gating and data association, conditions under which these assumptions may no longer hold.
Validation gating and NN association are standard components of tracking systems in aerospace, radar, and navigation applications [3, 17, 19, 11]. Despite their widespread use, the statistical impact of these selection mechanisms on innovation-based diagnostics has received comparatively little explicit theoretical treatment.
While truncation of Gaussian distributions is classical, the implications of mandatory truncation and order-statistic selection on innovation-based consistency diagnostics have not been explicitly characterized in the Kalman tracking literature.
In particular, existing analyses typically focus on detection performance or association accuracy, rather than on the induced bias in innovation statistics themselves.
Recent work on adaptive and learning enhanced Kalman filtering further highlights the reliance on innovation statistics for online consistency assessment and tuning [13, 14, 15]. These methods benefit from a precise understanding of how structural elements of the tracking pipeline, such as gating and association, affect the observed innovation process.
The present work complements the existing literature by providing an explicit statistical characterization of gate-conditioned innovation moments and by showing that NN association introduces an unavoidable order-statistic bias. Unlike prior adaptive or robust filtering approaches, the results here isolate selection-induced effects under ideal modeling assumptions, thereby clarifying fundamental limitations in the interpretation of innovation-based diagnostics.
The main contributions of this paper are summarized as follows:
  • Validation gating is interpreted as a statistical conditioning operation, and exact expressions are derived for the first- and second-order moments of the gate-conditioned innovation under linear–Gaussian assumptions.
  • It is shown that ellipsoidal gating induces a deterministic contraction of the innovation covariance that depends only on the gate threshold and the measurement dimension.
  • NN association is characterized as an order-statistic selection mechanism, and it is proven that this selection introduces an additional, unavoidable contraction of innovation energy, leading to an impossibility result: nominal innovation statistics cannot be preserved under nontrivial gating and association.
  • Closed-form results and illustrative examples are provided for the two-dimensional case, quantifying the combined effects of gating and association and highlighting their practical significance.
The remainder of the paper is organized as follows. Section II introduces the problem formulation and the innovation model. Section III analyzes validation gating as a statistical selection mechanism and derives gate-conditioned innovation moments. Section IV discusses the implications of these results for innovation-based consistency diagnostics. Section V analyzes the additional bias induced by NN association. Section VI presents a two-dimensional case study with closed-form expressions and illustrative examples. Section VII concludes the paper and discusses implications and limitations.
II Problem Formulation
II-A Kalman Filter Innovation Model
The analysis is conducted within the classical linear Gaussian Kalman filtering framework.
Consider the discrete-time linear measurement model
$z_{k}=Hx_{k}+v_{k},$ (1)
where $x_{k}\in\mathbb{R}^{n}$ denotes the system state, $z_{k}\in\mathbb{R}^{m}$ the measurement vector, $H\in\mathbb{R}^{m\times n}$ the measurement matrix, and $v_{k}\sim\mathcal{N}(0,R)$ the measurement noise.
Let $\hat{x}_{k}^{-}$ and $P_{k}^{-}$ denote the predicted state estimate and error covariance produced by a Kalman filter. The innovation is defined as
$\nu_{k}\triangleq z_{k}-H\hat{x}_{k}^{-},$ (2)
with associated innovation covariance
$S_{k}\triangleq HP_{k}^{-}H^{\top}+R.$ (3)
Under nominal Kalman filter assumptions, the innovation sequence $\{\nu_{k}\}$ is zero-mean, Gaussian, and white. In the sequel, we analyze a single time index and omit the subscript $k$ for notational clarity. Accordingly, we model the innovation as the random vector
$\nu:\Omega\rightarrow\mathbb{R}^{m},\qquad\nu\sim\mathcal{N}(0,S),$ (4)
where $S\in\mathbb{S}_{++}^{m}$ denotes the symmetric positive-definite innovation covariance matrix.
II-B Post-Gate Innovation Distribution
Under the nominal linear Gaussian Kalman filter assumptions, the normalized innovation squared (NIS) is defined as
$Z\triangleq\nu^{\top}S^{-1}\nu,$ (5)
and follows a chi-square distribution with $m$ degrees of freedom.
Let
$\mathcal{A}\triangleq\{Z\leq\tau\}$ (6)
denote the validation (acceptance) event induced by ellipsoidal gating. All innovation samples that pass the gate are therefore distributed according to the conditional distribution of $\nu$ given $\mathcal{A}$, denoted by $\nu\mid\mathcal{A}$. This conditional distribution corresponds to a Gaussian law truncated to the ellipsoidal region
$\mathcal{E}_{\tau}\triangleq\{\nu\in\mathbb{R}^{m}:\mathcal{A}\}$
rather than the unconditional Gaussian distribution in (4). The statistical properties of this gate-conditioned innovation distribution form the basis of the analysis in the remainder of this paper.
III Innovation Gating as a Statistical Selection Mechanism
We interpret ellipsoidal gating as a statistical selection mechanism acting on the innovation process. Specifically, gating restricts the observed innovation stream to realizations satisfying the acceptance event $\mathcal{A}$ defined in (6). Consequently, all innovations that enter data association, state update, and diagnostic logic are drawn from the conditional distribution $\nu\mid\mathcal{A}$, rather than from the nominal Gaussian distribution $\mathcal{N}(0,S)$.
III-A Distribution of the NIS and Ellipsoidal Gating
Proposition 1 (Distribution of the NIS).
If $\nu\sim\mathcal{N}(0,S)$ with $S\in\mathbb{S}_{++}^{m}$, then
$Z\sim\chi^{2}_{m}.$ (7)
This classical result provides the probabilistic basis for ellipsoidal validation gating: selecting the threshold $\tau$ as a chi-square quantile ensures
$\mathbb{P}\{\mathcal{A}\}=P_{g}$ (8)
under the nominal innovation model.
Once validation gating is applied, however, the innovation process is no longer observed unconditionally, but only through realizations satisfying the acceptance event $\mathcal{A}=\{Z\leq\tau\}$. As a result, all innovation-based statistics computed after gating are statistics of a conditionally observed random variable. In the innovation space $\mathbb{R}^{m}$, the acceptance event $\mathcal{A}$ corresponds to the ellipsoidal region $\mathcal{E}_{\tau}$. Validation gating therefore implements a deterministic truncation of the innovation distribution to $\mathcal{E}_{\tau}$, retaining only realizations within a fixed Mahalanobis radius.
III-B Gate-Conditioned Innovation Moments
Proposition 2 (Gate-Conditioned Innovation Moments).
Let $\nu\sim\mathcal{N}(0,S)$ with $S\in\mathbb{S}_{++}^{m}$, and let $\mathcal{A}=\{\nu^{\top}S^{-1}\nu\leq\tau\}$ denote the validation event. Then the gate-conditioned innovation satisfies
$\mathbb{E}[\nu\mid\mathcal{A}]$ $=0,$ (9a)
$\mathbb{E}[\nu\nu^{\top}\mid\mathcal{A}]$ $=\gamma(\tau,m)\,S,$ (9b)
where the scalar contraction factor $\gamma(\tau,m)$ is given by
$\gamma(\tau,m)\triangleq\frac{1}{m}\,\mathbb{E}\!\left[Z\mid Z\leq\tau\right],\qquad Z\sim\chi^{2}_{m},$ (10)
and satisfies $0<\gamma(\tau,m)<1$.
Proposition 2 shows that ellipsoidal gating induces a deterministic, dimension-dependent contraction of the innovation covariance, while preserving zero mean. Importantly, the contraction factor depends only on the gate threshold $\tau$ and the measurement dimension $m$, and is independent of the nominal covariance matrix $S$.
III-C Gate-Conditioned NIS Statistics
An immediate consequence of (9b) is an explicit expression for the mean NIS after gating.
Corollary 1 (Gate-Conditioned Mean NIS).
Under the assumptions of Proposition 2,
$\mathbb{E}[Z\mid\mathcal{A}]=m\,\gamma(\tau,m),$ (11)
where $Z=\nu^{\top}S^{-1}\nu$.
In contrast to the nominal reference value $\mathbb{E}[Z]=m$, the expected NIS computed from accepted measurements reflects the conditioning induced by the validation gate. Fig. 1 illustrates the deterministic contraction of NIS statistics induced by validation gating in the two-dimensional case.
Figure 1: Gate-conditioned NIS distribution in the two-dimensional case. The nominal $\chi^{2}_{2}$ distribution is shown together with the truncated distribution induced by ellipsoidal validation gating. The gate-conditioned mean is systematically lower than the nominal reference value, illustrating the systematic contraction of innovation energy relative to the nominal reference induced by validation gating, even under ideal Kalman filter assumptions.
III-D Implications for Innovation-Based Diagnostics
Once ellipsoidal gating is applied, innovation-based statistics computed from accepted measurements converge to conditional, rather than nominal, quantities. In particular,
$\mathbb{E}[\nu\nu^{\top}\mid\mathcal{A}]\neq S,\qquad\mathbb{E}[Z\mid\mathcal{A}]\neq m.$ (12)
These deviations arise solely from the conditioning induced by validation gating and persist even under perfectly matched system and noise models. Consequently, post-gate innovation covariance estimates and NIS statistics cannot be interpreted using unconditional Gaussian or chi-square reference values.
Any diagnostic or adaptive tuning procedure operating on post-gate innovation statistics is therefore estimating gate-conditioned quantities. Interpreting post-selection innovation statistics using nominal, unconditional reference values systematically leads to the appearance of reduced innovation energy relative to the nominal covariance.
When nearest-neighbor association is applied, this effect is further amplified by order-statistic selection and becomes pronounced in high-density scenarios, precisely in regimes where NN association is known to be theoretically suboptimal.
IV Gate-Aware Consistency and Interpretation of Innovation Statistics
The results of Section III establish that ellipsoidal gating alters the statistical properties of the innovation process. In this section, we examine the implications of these results for innovation-based consistency diagnostics, with particular emphasis on NIS.
IV-A Classical use of NIS for consistency Validation
In Kalman-based tracking systems, the normalized innovation $Z_{k}$ is commonly employed as a diagnostic quantity for filter consistency.
Under the nominal Kalman filter assumptions recalled in Section III, the NIS follows a chi-square distribution with $m$ degrees of freedom. As a result, standard practice is to compare instantaneous NIS values to chi-square confidence bounds, or to monitor the empirical mean of $\{Z_{k}\}$ over time and compare it to the nominal reference value $m$, as commonly done in online consistency testing [16].
If the empirical mean is significantly below $m$, the filter is often declared overconfident; if it is significantly above $m$, it is declared underconfident.
IV-B Post-gate NIS is not chi-square
As a direct consequence of the conditioning induced by ellipsoidal gating, the normalized innovation squared no longer follows its nominal distribution. Specifically,
$Z\mid\mathcal{A}\;\not\sim\;\chi^{2}_{m},$ (13)
but instead follows a truncated chi-square law obtained by renormalizing $\chi^{2}_{m}$ on the interval $0,\tau] [$6].
Accordingly, as established in Proposition 2,
$\mathbb{E}[Z\mid\mathcal{A}]=m\,\gamma(\tau,m),$ (14)
with $\gamma(\tau,m)\in(0,1)$. The nominal reference value $\mathbb{E}[Z]=m$ therefore ceases to be valid whenever gating is applied.
IV-C Gate-aware interpretation of NIS statistics
The results above imply that standard NIS diagnostics must be interpreted relative to gate-conditioned reference values. Specifically, when NIS statistics are computed from accepted innovations only, the correct reference mean is given in Proposition 2. Equivalently, a gate-aware normalized statistic is defined as
$Z_{k}^{\text{corr}}\triangleq\frac{1}{\gamma(\tau,m)}\,Z_{k},$ (15)
which satisfies
$\mathbb{E}[Z_{k}^{\text{corr}}\mid\mathcal{A}]=m.$ (16)
This normalization allows classical chi-square intuition to be reused without modifying the underlying Kalman filter or gating logic, while preserving consistency interpretations used in practice [16].
IV-D Implications for adaptive tuning and diagnostics
Many adaptive noise-tuning and consistency-monitoring schemes rely, either explicitly or implicitly, on innovation covariance estimates or NIS-based statistics [7], and often enforce nominal chi-square innovation behavior as a tuning objective [5].
If such schemes operate on post-gate data without accounting for the conditioning induced by gating, they will systematically underestimate the innovation covariance, reduce estimated measurement noise levels, tighten validation gates, and increase the probability of missed detections.
This feedback mechanism arises even under ideal modeling assumptions and is a direct consequence of ignoring the gate-conditioned nature of the observed innovation process.
The contraction factor $\gamma(\tau,m)$ admits intuitive limiting behavior. As $\tau\to\infty$, validation gating becomes inactive and $\gamma(\tau,m)\to 1$, recovering the nominal innovation statistics. Conversely, as $\tau\to 0$, only vanishingly small innovations are accepted and $\gamma(\tau,m)\to 0$, driving the post-gate innovation energy to zero.
V Bias Induced by Nearest-Neighbor Association
The previous sections characterized the statistical effect of ellipsoidal gating on innovation moments. In practical tracking systems, however, gating is only the first stage of a selection process. When multiple measurements fall inside the validation region, a data association rule is applied to select a single measurement for the state update. The most widely used rule in classical tracking is NN association [18, 2].
This section shows that NN association introduces an additional and unavoidable statistical bias that compounds the gate-induced effects. Unlike modeling errors or tuning artifacts, this bias arises solely from order-statistic selection and persists even under ideal nominal assumptions.
V-A Nearest-neighbor association as statistical selection
Consider a time step at which $M\geq 1$ measurements pass the validation gate $\mathcal{A}$. Let $\{\nu^{(i)}\}_{i=1}^{M}$ denote the corresponding innovation vectors, each generated according to the same post-gate distribution $(\nu\mid\mathcal{A})$. NN association selects the innovation with minimum NIS, as originally proposed in the classical multi-target tracking framework of Reid [18],
$i^{\star}\triangleq\arg\min_{1\leq i\leq M}\|\nu^{(i)}\|_{S^{-1}}^{2},$ (17)
and uses $\nu^{(i^{\star})}$ for the update step.
While (17) is often motivated algorithmically as a simple and efficient approximation to more complex association schemes, it constitutes a nonlinear statistical selection operator acting directly on the innovation.
V-B Energy contraction under NN association
The central effect of NN association is most clearly exposed at the level of second-order energy. The following proposition formalizes this effect without invoking Gaussianity or any specific parametric form.
Proposition 3 (Post-gate NN energy contraction).
Let $\nu\in\mathbb{R}^{d}$ be an innovation vector and let $\mathcal{A}$ denote the gating acceptance event. Let $\{\nu^{(i)}\}_{i=1}^{M}$ be independent samples from the conditional distribution $(\nu\mid\mathcal{A})$, and let $i^{\star}=\arg\min_{i}\|\nu^{(i)}\|$. Then
$\mathbb{E}\!\left[\|\nu^{(i^{\star})}\|^{2}\mid\mathcal{A}\right]\;\leq\;\mathbb{E}\!\left[\|\nu\|^{2}\mid\mathcal{A}\right],$ (18)
with strict inequality for $M>1$ whenever the conditional distribution $(\nu\mid\mathcal{A})$ is non-degenerate.
This bound shows that NN association induces a systematic contraction of innovation energy beyond that caused by gating alone. The effect is purely statistical and follows from order-statistic selection.
Corollary 2 (Impossibility of preserving nominal innovation energy under selection).
Let $\nu\sim\mathcal{N}(0,S)$ with $S\in\mathbb{S}_{++}^{m}$ and let $\mathcal{A}=\{\nu^{\top}S^{-1}\nu\leq\tau\}$ be a nontrivial ellipsoidal gate with $\Pr(\mathcal{A})\in(0,1)$. Assume that, whenever the gate admits $M\geq 1$ candidate measurements, the association rule selects the minimum-norm innovation among $M$ independent post-gate candidates (NN association), yielding $\nu^{(i^{\star})}$.
Then, for any $M>1$ for which $(\nu\mid\mathcal{A})$ is non-degenerate, the selected (post-gate, post-association) innovation cannot preserve the nominal innovation energy:
$\mathbb{E}\!\left[\|\nu^{(i^{\star})}\|^{2}\right]\;<\;\mathbb{E}\!\left[\|\nu\|^{2}\right]\;=\;\operatorname{tr}(S).$ (19)
In particular, no choice of the gate threshold $\tau$ with $\Pr(\mathcal{A})\in(0,1)$ can make the post-selection innovation energy match its nominal value when $M>1$ occurs with nonzero probability.
Proof.
By the law of total expectation,
$\mathbb{E}\!\left[\|\nu^{(i^{\star})}\|^{2}\right]$ $=\Pr(\mathcal{A})\,\mathbb{E}\!\left[\|\nu^{(i^{\star})}\|^{2}\mid\mathcal{A}\right]$
$\quad+\Pr(\mathcal{A}^{c})\,\mathbb{E}\!\left[\|\nu^{(i^{\star})}\|^{2}\mid\mathcal{A}^{c}\right].$
Since $\nu^{(i^{\star})}$ is defined only when the gate accepts, we interpret the post-selection innovation stream as the accepted stream, so that the relevant energy is $\mathbb{E}[\|\nu^{(i^{\star})}\|^{2}\mid\mathcal{A}]$. Proposition 5.1 gives, for $M>1$ and non-degenerate $(\nu\mid\mathcal{A})$,
$\mathbb{E}\!\left[\|\nu^{(i^{\star})}\|^{2}\mid\mathcal{A}\right]<\mathbb{E}\!\left[\|\nu\|^{2}\mid\mathcal{A}\right].$
Moreover, Proposition 2 implies
$\mathbb{E}\!\left[\|\nu\|^{2}\mid\mathcal{A}\right]$ $=\operatorname{tr}\!\left(\mathbb{E}[\nu\nu^{\top}\mid\mathcal{A}]\right)$
$=\operatorname{tr}\!\left(\gamma(\tau,m)\,S\right)=\gamma(\tau,m)\,\operatorname{tr}(S).$
with $\gamma(\tau,m)\in(0,1)$ for any nontrivial gate $\Pr(\mathcal{A})\in(0,1)$. Therefore,
$\mathbb{E}\!\left[\|\nu^{(i^{\star})}\|^{2}\mid\mathcal{A}\right]<\gamma(\tau,m)\,\operatorname{tr}(S)<\operatorname{tr}(S)=\mathbb{E}\!\left[\|\nu\|^{2}\right],$
which yields (19) and proves the claim. ∎
V-C Interpretation via order statistics
The purpose of this subsection is to provide intuition for Proposition 5.1 by isolating the purely statistical mechanism underlying NN association. The mechanism underlying (18) is elementary. NN association selects the minimum-norm element from a finite set of independent post-gate innovations. As a result, the selected innovation corresponds to the first-order statistic of the squared norms.
For any nonnegative, non-degenerate random variable $X$ and $M\geq 2$ independent copies $\{X^{(i)}\}$,
$\mathbb{E}\!\left[\min_{i}X^{(i)}\right]<\mathbb{E}[X].$ (20)
Applying this property to $X=\|\nu\|^{2}\mid\mathcal{A}$ yields the strict inequality in (18). Notably, this argument relies only on elementary properties of order statistics and does not invoke Gaussianity, dimensionality, or the specific shape of the validation gate.
Fig. 2 visualizes the order-statistic bias introduced by NN association as the number of in-gate measurements increases.
Figure 2: Effect of NN association on normalized innovation squared (NIS) statistics after gating. The distribution of $Z\mid\mathcal{A}$ is shown together with the distribution of the minimum NIS selected from $M=2,3,5$ independent in-gate measurements. As $M$ increases, the expected selected NIS decreases, illustrating the order-statistic energy contraction induced by nearest-neighbor association.
V-D Compound selection pipeline
Combining gating and NN association, the effective selection pipeline acting on the innovation can be summarized as
$\nu\;\xrightarrow{\text{gating}}\;\nu\mid\mathcal{A}\;\xrightarrow{\text{NN association}}\;\nu^{(i^{\star})}.$ (21)
Gating induces a deterministic contraction of innovation energy through conditioning, as characterized in previous sections. NN association then applies an additional contraction through order-statistic selection, as quantified by (18).
At the level of covariance structure, this compound effect can be summarized conceptually as a further shrinkage relative to the gate-conditioned covariance.
V-E Implications for innovation-based diagnostics
The bound (18) has direct implications for innovation-based diagnostics and tuning. Statistics computed from NN-selected innovations exhibit systematically reduced energy relative to both nominal and gate-conditioned references. Consequently, diagnostic rules based on expected NIS values will indicate a systematic reduction of innovation energy relative to nominal reference values, even when the system operates exactly according to the nominal model. Crucially, this behavior does not arise from clutter, false alarms, or modeling errors. It is a structural consequence of NN association itself. As such, it cannot be eliminated by tuning and must be explicitly accounted for when interpreting innovation statistics in gated tracking systems.
VI Two-Dimensional Tracking Case Study
This section illustrates the theoretical results of Sections III–V in a two-dimensional tracking scenario. The case $m=2$ is of particular practical relevance for planar position measurements and admits closed-form expressions that make the effects of leading selection mechanisms quantitatively explicit.
VI-A State-space and innovation model
We consider a standard linear state-space model for planar target tracking. The state vector
$x_{k}=\begin{bmatrix}p_{x,k}&p_{y,k}&v_{x,k}&v_{y,k}\end{bmatrix}^{\top}\in\mathbb{R}^{4}$ (22)
evolves according to a constant-velocity model,
$x_{k}=Fx_{k-1}+w_{k-1},$ (23)
where $w_{k}$ is zero-mean Gaussian process noise. The measurement model observes position only,
$z_{k}=Hx_{k}+v_{k},$ (24)
with zero-mean Gaussian measurement noise $v_{k}$.
Under nominal Kalman filter assumptions, the innovation $\nu_{k}=z_{k}-H\hat{x}_{k|k-1}$ is zero-mean Gaussian with covariance $S_{k}\in\mathbb{R}^{2\times 2}$. In the following analysis, a fixed time index is considered and the subscript $k$ is omitted for clarity.
VI-B Innovation statistics in two dimensions
For $m=2$, the NIS, $Z=\nu^{\top}S^{-1}\nu$, follows a chi-square distribution with two degrees of freedom under nominal linear Gaussian assumptions. Equivalently, $Z$ is exponentially distributed with density
$f_{Z}(z)=\frac{1}{2}e^{-z/2},\qquad z\geq 0,$ (25)
and cumulative distribution function
$F_{Z}(z)=1-e^{-z/2}.$ (26)
Given a validation gate with acceptance probability $P_{g}$, the gate threshold $\tau$ is defined implicitly by $\Pr(Z\leq\tau)=P_{g}$. In the two-dimensional case this yields the closed-form expression
$\tau=-2\ln(1-P_{g}).$ (27)
The post-gate innovation distribution is therefore obtained by conditioning the exponential law of $Z$ on the event $Z\leq\tau$, i.e., by truncation to the interval $0,\tau]$ and renormalization.
VI-C Explicit gate-induced contraction
As shown in Section IV, ellipsoidal gating induces a deterministic contraction of innovation energy. In two dimensions, this contraction admits a closed-form characterization due to the exponential structure of the NIS distribution.
The gate-conditioned mean NIS is given by
$\mathbb{E}[Z\mid Z\leq\tau]=\frac{\int_{0}^{\tau}zf_{Z}(z)\,dz}{\int_{0}^{\tau}f_{Z}(z)\,dz}=\frac{2-(\tau+2)e^{-\tau/2}}{1-e^{-\tau/2}}.$ (28)
Substituting $\tau=-2\ln(1-P_{g})$ and simplifying yields
$\mathbb{E}[Z\mid Z\leq\tau]=2\left(1+\frac{(1-P_{g})\ln(1-P_{g})}{P_{g}}\right).$ (29)
Using the relation $\mathbb{E}[\nu\nu^{\top}\mid\mathcal{A}]=\frac{1}{m}\mathbb{E}[Z\mid\mathcal{A}]\,S$ with $m=2$, the gate-conditioned innovation covariance can be written as
$\mathbb{E}[\nu\nu^{\top}\mid\mathcal{A}]=\gamma(P_{g},2)\,S,$ (30)
where the contraction factor is
$\gamma(P_{g},2)=1+\frac{(1-P_{g})\ln(1-P_{g})}{P_{g}}.$ (31)
Table I reports $\gamma(P_{g},2)$ and the corresponding post-gate mean $\mathbb{E}[Z\mid\mathcal{A}]$ for several commonly used gate probabilities. Even for moderate values of $P_{g}$, the effective innovation energy is significantly reduced relative to the nominal model.
Gate probability $P_{g}$ Threshold $\tau$ $\gamma(P_{g},2)$ $\mathbb{E}[Z\mid\mathcal{A}]$
0.90 4.605 0.744 1.488
0.95 5.991 0.842 1.684
0.99 9.210 0.953 1.906
TABLE I: Gate-induced contraction and post-gate NIS mean in the two-dimensional case.
VI-D Interpretation
The two-dimensional case provides a concrete numerical instantiation of the general results derived earlier. In this setting, the effect of ellipsoidal gating can be quantified explicitly, revealing a predictable contraction of innovation energy that depends only on the gate probability. NN association further compounds this effect through order statistic selection. As illustrated in Table I, even commonly used gate probabilities lead to a substantial reduction in post-gate innovation energy. Consequently, innovation-based diagnostics computed after gating or NN association, when interpreted using nominal reference values, systematically indicate a reduction of innovation energy relative to the nominal reference.
VII Discussion
The analysis developed in this paper provides a principled explanation for a widely observed phenomenon in Kalman-based tracking systems: the tendency of innovation-based diagnostics evaluated after gating and association to exhibit systematically reduced innovation energy relative to nominal reference values, even in well-functioning trackers.
Sections 3-6 show that this behavior arises from two structural selection mechanisms inherent to practical tracking pipelines:
  • Ellipsoidal gating, which deterministically truncates the innovation distribution and contracts its covariance, and
  • NN association, which introduces an additional order-statistic bias by selecting the minimum-norm innovation among multiple candidates.
Both mechanisms operate even under ideal modeling assumptions. They do not rely on clutter, false alarms, nonlinear dynamics, or model mismatch. As a result, tuning-based explanations alone are insufficient to account for systematic bias observed in post-update innovation statistics.
VII-A Implications for Classical Tracking Theory
The apparent discrepancy between nominal and observed innovation statistics arises because classical theory characterizes the unconditional innovation process, whereas practical tracking systems operate on a conditionally observed innovation stream.
We do not propose an alternative to adaptive or robust Kalman filtering. Rather, providing a missing statistical layer that clarifies how innovation-based diagnostics should be interpreted prior to invoking adaptation mechanisms. In this sense, the analysis complements existing tuning and robustness methods by separating structural selection effects from genuine model mismatch.
NN association is treated here as a statistical selection operator whose impact on innovation statistics can be characterized independently of clutter models or hypothesis management.
VII-B Implications for implementation and evaluation
From an implementation perspective, the results suggest several practical guidelines:
  • Innovation covariance and NIS statistics computed after gating should be interpreted using gate-conditioned reference values.
  • Systematically reduced innovation energy observed in post-gate diagnostics does not necessarily indicate filter mis-tuning.
  • Adaptive tuning schemes based on innovation statistics should explicitly account for selection-induced bias to avoid self-reinforcing gate tightening.
These guidelines can be incorporated into existing tracking systems without modifying the Kalman recursion, validation logic, or association rules themselves.
VII-C Limitations and scope
The analysis in this paper is deliberately restricted to linear measurement models, Gaussian noise, ellipsoidal validation gates, and NN association. These restrictions ensure that the identified effects are not artifacts of nonlinearity or non-Gaussianity.
Extensions to more complex settings, such as nonlinear filters, probabilistic data association, or explicit clutter models, are natural directions for future work.
VIII Conclusion
Validation gating and NN association introduce predictable and quantifiable biases in innovation statistics, even under ideal Kalman filter assumptions. Ellipsoidal gating deterministically contracts the innovation covariance, while NN association introduces an additional multiplicity- dependent bias through order-statistic selection.
These structural effects explain why innovation-based diagnostics evaluated after gating and association systematically deviate from nominal chi-square references in practical tracking systems.
Acknowledgment
The author gratefully acknowledges Prof. Yaakov Bar-Shalom for his insightful comments that helped clarify the interpretation and terminology of the paper.
Appendix A Proof of Proposition 2
We prove the statements in Proposition 2 by transforming the innovation into whitened coordinates and exploiting rotational invariance of the Gaussian distribution.
A-A Whitening Transformation
Since $S\in\mathbb{S}_{++}^{m}$, there exists a symmetric matrix $S^{1/2}$ such that $S^{1/2}S^{1/2}=S$. Define the whitened innovation
$u\triangleq S^{-1/2}\nu.$ (32)
By construction,
$u\sim\mathcal{N}(0,I_{m}),$ (33)
and the NIS becomes
$Z=\nu^{\top}S^{-1}\nu=u^{\top}u=\|u\|^{2}.$ (34)
Accordingly, the gating event can be written as
$\mathcal{A}=\{\|u\|^{2}\leq\tau\}.$ (35)
A-B Conditional Mean
The distribution of $u$ is symmetric about the origin, and the event $\|u\|^{2}\leq\tau$ is invariant under the transformation $u\mapsto-u$. Therefore,
$\mathbb{E}[u\mid\mathcal{A}]=0.$ (36)
Transforming back to the original coordinates yields
$\mathbb{E}[\nu\mid\mathcal{A}]=S^{1/2}\mathbb{E}[u\mid\mathcal{A}]=0,$ (37)
which establishes (9a).
A-C Conditional Covariance
Define the conditional second-order moment of the whitened innovation as
$C\triangleq\mathbb{E}[uu^{\top}\mid\mathcal{A}].$ (38)
Since $u\sim\mathcal{N}(0,I_{m})$ is rotationally invariant and the event $\|u\|^{2}\leq\tau$ depends only on the norm of $u$, the conditional distribution $u\mid\mathcal{A}$ is also rotationally invariant. Consequently, for any orthogonal matrix $Q\in\mathbb{R}^{m\times m}$,
$Qu\mid\mathcal{A}\;\overset{d}{=}\;u\mid\mathcal{A},$ (39)
which implies
$QCQ^{\top}=C.$ (40)
The only matrices commuting with all orthogonal transformations are scalar multiples of the identity. Therefore, there exists a scalar $\gamma(\tau,m)>0$ such that
$C=\gamma(\tau,m)\,I_{m}.$ (41)
Taking the trace of (41) gives
$\mathrm{tr}(C)=m\,\gamma(\tau,m).$ (42)
On the other hand,
$\mathrm{tr}(C)=\mathbb{E}[\mathrm{tr}(uu^{\top})\mid\mathcal{A}]=\mathbb{E}[u^{\top}u\mid\mathcal{A}]=\mathbb{E}[Z\mid Z\leq\tau].$ (43)
Combining the above expressions yields
$\gamma(\tau,m)=\frac{1}{m}\,\mathbb{E}[Z\mid Z\leq\tau],$ (44)
which establishes (10).
Finally, transforming back to the original coordinates,
$\mathbb{E}[\nu\nu^{\top}\mid\mathcal{A}]=S^{1/2}CS^{1/2}=\gamma(\tau,m)\,S,$ (45)
which proves (9b).
A-D Bounds on $\gamma(\tau,m)$
Since $Z\sim\chi^{2}_{m}$ satisfies $\mathbb{E}[Z]=m$ and the event $Z\leq\tau$ truncates the right tail of the distribution,
$0<\mathbb{E}[Z\mid Z\leq\tau]<m.$ (46)
Dividing by $m$ yields
$0<\gamma(\tau,m)<1,$ (47)
which completes the proof.
Appendix B Proof of Proposition 5.1
We provide the proof of the post-gate NN energy contraction bound.
Proof.
Condition on the gating acceptance event $\mathcal{A}$. Let $\nu^{(1)},\dots,\nu^{(M)}$ be independent samples from the conditional distribution $(\nu\mid\mathcal{A})$, and define the nonnegative random variables
$X_{i}\;\triangleq\;\|\nu^{(i)}\|^{2},\qquad i=1,\dots,M,$
and
$X\;\triangleq\;\|\nu\|^{2}\quad\text{under the same conditional law }(\nu\mid\mathcal{A}).$
By construction, $\{X_{i}\}_{i=1}^{M}$ are i.i.d. and each $X_{i}$ has the same distribution as $X$ given $\mathcal{A}$.
NN association selects the minimum-norm innovation, hence
$X_{(1)}\;\triangleq\;\|\nu^{(i^{\star})}\|^{2}\;=\;\min_{1\leq i\leq M}X_{i}.$
Since $X_{(1)}\leq X_{1}$ almost surely, taking conditional expectations yields
$\mathbb{E}\!\left[X_{(1)}\mid\mathcal{A}\right]\;\leq\;\mathbb{E}\!\left[X_{1}\mid\mathcal{A}\right]\;=\;\mathbb{E}\!\left[\|\nu\|^{2}\mid\mathcal{A}\right],$
which establishes the inequality in Proposition 5.1.
For strict inequality when $M>1$, assume that the conditional distribution of $X$ is non-degenerate. Then there exists a set of positive probability on which $X_{2}<X_{1}$, implying
$\Pr\!\left(X_{(1)}<X_{1}\mid\mathcal{A}\right)>0.$
Consequently,
$\mathbb{E}\!\left[X_{(1)}\mid\mathcal{A}\right]<\mathbb{E}\!\left[X_{1}\mid\mathcal{A}\right],$
which completes the proof. ∎